OpenStudy (anonymous):

Find the equation of the ellipse with vertices (0, -1) and (12, -1) and minor axis of length 6 NEED HELP A.S.A.P

OpenStudy (amistre64):

the general setup to start with will be:\[(\frac{(x-h)}{a})^2+(\frac{(y-k)}{b})^2=1\]for a center of (h,k)

OpenStudy (amistre64):



OpenStudy (amistre64):

looks like b is 3 and a is 6 to me

OpenStudy (anonymous):

THANK-YOU SO VERY MUCH!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!